OpenStudy (anonymous):

A student planning her curriculum for the upcoming year must select one of four business courses, one of three mathematics courses, two of eight elective courses, and either one of two history courses or one of three social science courses. How many different curricula are available for her consideration?

OpenStudy (anonymous):

did you have a question about my answer last time? you take all the choices and multiply them together

OpenStudy (anonymous):

\[4\times 3\times 28\times 5\]

OpenStudy (anonymous):

oh wow i thought i did it right now i get it lol thanks :)

OpenStudy (anonymous):

yw